OpenStudy (unklerhaukus):

xy'+y+x^{2}y^{2}e^{x}=0

OpenStudy (unklerhaukus):

\[xy'+y+x^{2}y^{2}e^{x}=0\]

OpenStudy (anonymous):

Bernoulli Equation \[x\frac{dy}{dx}+e^{x}x^{2}y^{2}+y=0\]\[x\frac{dy}{dx}+y=-e^{x}x^{2}y^{2}\]\[-\frac{1}{xy^{2}}x\frac{dy}{dx}-\frac{1}{xy}=e^{x}\] \[v=\frac{1}{y}\]\[\frac{dv}{dx}=-\frac{1}{y^2}\frac{dy}{dx}\] \[\frac{dv}{dx}=\frac{v}{x}=e^{x}x\] \[\mu=e^{\int-\frac{1}{x}dx}=\frac{1}{x}\] \[\frac{1}{x}\frac{dv}{dx}-\frac{v}{x^2}=e^{x}\] \[-\frac{1}{x^2}=\frac{d}{dx}\left(\frac{1}{x}\right)\] \[\frac{1}{x}\frac{dv}{dx}+\frac{d}{dx}\left(\frac{1}{x}\right)v=e^{x}\]\[\frac{d}{dx\left(\frac{v}{x}\right)}=e^{x}\]\[\frac{v}{x}=e^{x}+c_1\]\[v=x(e^{x}+c_1)\]\[y=\frac{1}{x(e^{x}+c_1)}\]pz

OpenStudy (unklerhaukus):

i knew it was non linear , thankyou

OpenStudy (unklerhaukus):

im confused on your third line did you divide the second line by \[-1/(x^{2}y^{2})\]